OpenStudy (anonymous):

[Reward:Medal & Fan] If the opposite angles of a parallelogram are (2x-3)DEGREE and (45-x)degree,then the value of x is? Please Help!!!

OpenStudy (anonymous):

@wio help!

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

what do we know about opposite angles of a parallelogram?

OpenStudy (anonymous):

@zzr0ck3r They are equal?

OpenStudy (zzr0ck3r):

so we have \(2x-3=45-x\)

OpenStudy (zzr0ck3r):

Can you solve now?

OpenStudy (anonymous):

@zzr0ck3r No plz

OpenStudy (zzr0ck3r):

add x to both sides and what do we have?

OpenStudy (zzr0ck3r):

\(2x-3=45-x\iff3x-3=45\iff3x=48\iff x=16\)

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

np